Circular No. 9054

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2468 CYGNI M. L. Sitko, Space Science Institute; and D. K. Lynch, The Aerospace

Corporation, report on 0.8- to 2.4-μm spectroscopy of V2468 Cyg (cf. IAUCs 8936, 8998) on June 18 UT using SpeX at the Infrared Telescope Facility. The object continues to develop in its coronal phase, showing lines of [Si VIII] 0.9913-μm, [Fe VI] 1.2330-μm, [Fe VI] 1.2674-μm, [Si VI] 1.964μm, [Al IX] 2.0444-μm, [Ca VIII] 2.32-μm, and possibly [Ti VI] 1.7155-μm and [P VIII] 1.7356-μm. Many He II lines are also present. Rydberg lines at 0.8926, 1.1114, 1.5545, and 2.103 μm are also present (Lynch et al. 2008, Bull. AAS 40, 531). These lines generally have FWHM ∼ 2000 km/s, and there is no indication of dust formation.

COMETS C/2009 B6, 2009 B7, 2009 C1, AND 2009 C2 (STEREO) Further to IAUCs 9053, additional small and slightly diffuse Kreutz

sungrazers have been found on STEREO/SECCHI HI-1A images. Very approximate peak magnitudes from K. Battams: C/2009 B6 and C/2009 B7, 11; C/2009 C1, 11-12; C/2009 C2, 10.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 B6 Jan. 19.683 23^{h}21.^{m}2 − 8^{o}10′ HI* RM 2009-M37
C/2009 B7  29.739 23 57.1 − 3 43 HI* RM 2009-M37
C/2009 C1 Feb. 10.044 0 37.9 + 0 46 HI* RK 2009-M37
C/2009 C2  11.822 0 51.3 + 1 05 HI* KB 2009-M37

(3200) PHAETHON K. Battams, Naval Research Laboratory, also writes that A. Wat-

son (Werribee, Victoria, Australia) has commented that the minor planet (3200) was visible in SECCHI HI-1A images during June 17–22, noting a very short radial elongation (perpendicular to the direction of motion) that was possibly a line-of-sight effect related to its passage through a reasonably dense, higher-speed solar outflow stream. Battams adds that the apparent brightness of (3200) increased significantly (∼ 2 mag or more), peaking at mag perhaps 10–11 a few hours after perihelion (T = June 20.302 TT, q = 0.140 AU); 36 hr later, the object’s had faded to magnitude roughly 13–14. Phaethon was also visible in HI-1B images during June 21–22. More formal photometry will be performed later.

2009 June 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green